# Revealing Nanostructures in High-Entropy Alloys via Machine-Learning Accelerated Scalable Monte Carlo Simulation

### v1

@@ Abstract
Large-scale simulations are a powerful tool for understanding material behavior and designing new materials. However, the computational limitations of atomistic-scale simulations—stemming from constraints in computational power, algorithmic efficiency, and accuracy—prevent us from modeling large systems, even at ground state, let alone at finite temperatures, with density functional theory (DFT)-level precision. Here, we introduce an efficient and Scalable Monte Carlo (SMC) method that overcomes the parallelization bottlenecks inherent in conventional MC simulation algorithms, reducing the computational complexity of an MC sweep from $O(N^{2})$ to $O(N)$. Leveraging the large degree of parallelization inherent in our method, we present a GPU-accelerated implementation, SMC_GPU, which enables the simulation of atomistic systems exceeding one billion atoms while maintaining the accuracy of density functional theory (DFT). Using this unprecedented capability, we perform large-scale thermodynamic simulations to investigate the microstructure evolution in the FeCoNiAlTi and MoNbTaW high-entropy alloys (HEAs). Our results reveal a rich diversity of nanoscale phenomena, including short-range and long-range order, nanoparticles, and nanophases. The size, composition, and morphology of these nanostructures, which can comprise millions of atoms and thus present significant challenges for traditional methods, are analyzed with high-accuracy atomistic simulations for the first time, to the best of our knowledge. Our simulations produce thermodynamic properties and nanostructures that align well with available theoretical and experimental data. More intriguingly, our results reveal that the intricate nanoscale interplay of order and disorder in high-entropy alloys (HEAs) stems from the combined effects of chemical complexity and temperature, offering valuable guidance for alloy design. This work underscores the promising potential of large-scale MC simulation in exploring the largely uncharted territory of nanoscale order-disorder evolution in chemically complex materials.
@@ 1 Introduction
A central goal of computational materials science is to predict the physical properties of materials using only fundamental inputs such as atomic species and physical constants. A widely adopted strategy towards this goal is the direct solution of the quantum equations governing electron dynamics: the Schrödinger equation for non-relativistic cases and the Dirac equation for relativistic cases. However, the complexity of quantum many-body interactions necessitates approximations. For such a purpose, density functional theory (DFT) [1] has emerged as a particularly successful approach, which approximates intractable many-body effects through an exchange-correlation functional within a single-body framework. This approximation significantly enhances computational efficiency, enabling the routine prediction of ground-state physical properties for systems with several hundred atoms. However, for systems requiring hundreds to thousands of atoms to simulate, such as nanodefects, non-stoichiometric compounds, and complex solid-solution alloys, conventional DFT becomes prohibitively expensive due to its intrinsic $O(N^{3})$ scaling with system size. A further complication arises when considering finite-temperature effects [2]. Even for relatively small systems, rigorously accounting for temperature necessitates the evaluation of an enormous number of configurations, making DFT-based finite-temperature simulations computationally intractable. For example, a direct thermodynamic simulation of the CuZn alloy using a 250-atom supercell requires the calculation of 600,000 DFT energies [3], a task demanding some of the world’s largest supercomputers. Consequently, many compelling materials science problems remain beyond the reach of state-of-the-art (SOTA) first-principles methods.
@@
One common strategy to accelerate atomistic simulations is to replace computationally expensive electronic structure calculations with computationally efficient atomistic models. From the perspective of renormalization group theory [4, 5], the atomistic models can be interpreted as the effective model obtained by integrating out the electronic degrees of freedom, which substantially reduces the computational cost. Indeed, traditional interatomic potentials, such as the embedded atom model (EAM), are widely employed for large-scale thermodynamic simulations. However, a critical limitation of these methods lies in their reliance on hand-crafted features and extensive trial-and-error tuning by domain experts, which restricts their generalization capability and predictive accuracy. On the other hand, the advent of machine learning (ML) for atomistic systems [6, 7, 8, 9] has introduced a new paradigm to the construction of effective atomistic models [10, 11, 12, 13, 14]. Instead of being predefined, the parameters in ML atomistic models are automatically determined by training from high-quality first-principles datasets, therefore rendering it possible for the model to automatically capture complex interatomic interactions [15, 16, 17, 18]. Well-trained ML models are typically orders of magnitude faster than DFT methods, while can still retain their high accuracy [19, 20, 21]. This renders it possible to carry out high-accuracy atomistic simulations using supercells containing millions of atoms [22, 23, 24, 25], far exceeding the thousands-of-atoms limits of conventional DFT approaches.
@@
However, most machine learning potentials have been developed primarily to accelerate molecular dynamics (MD) simulations [22, 23, 7, 26, 27], with relatively limited applications in Monte Carlo (MC) simulations [28, 29, 16, 30, 29]. This is notable given that MC represents one of the two cornerstone methods in atomistic simulations, alongside MD [31]. A key challenge lies in the intrinsic sequential updating nature of widely used MC algorithms, such as the Metropolis algorithm, where the MC trials are generally attempted site by site, which hinders large-scale parallelization. This is in stark contrast to MD, where all the atoms are updated simultaneously in a single step. The sequential updating nature of MC quickly becomes the bottleneck as the system size grows, which undermines the efficiency advantage of ML atomistic models when integrating the two. On the other hand, for many interesting phenomena [32, 33, 34], such as order-disorder transitions in chemically complex materials, MC simulations remain the only viable approach due to their ability to efficiently sample the entire phase space. Consequently, developing highly scalable MC algorithms that overcome the limitations of sequential updating is of paramount importance for enabling the widespread integration of MC and ML atomistic models.
@@
To overcome the parallelization bottleneck inherent in conventional MC simulations, we introduce the Scalable Monte Carlo (SMC) method, which is a generalized checkerboard algorithm [35] that reduces the computational complexity of an MC sweep from $O(N^{2})$ to $O(N)$. Compared to the checkerboard algorithm variant [35] that only applies to Ising interactions, we demonstrate that the SMC method can be generalized to non-Ising interactions via the introduction of a local-interaction-zone of interatomic interactions, which renders it suitable for integration with machine learning atomistic models. We employ our method to study high-entropy alloys (HEAs) [36, 37], a class of chemically complex materials that have received significant attention due to their exceptional mechanical properties. These properties include overcoming the traditional strength-conductivity trade-off [38, 39, 40], attributed to phenomena such as chemical short-range order [41, 42, 43], nanoprecipitates [39], and nanophases [44]. Understanding these features necessitates large-scale MC simulations. Beyond mechanical properties, the size and morphology of nanostructures in HEAs also offer promising opportunities for catalysis, sparking widespread interest in recent years [45, 46, 47]. We present a GPU-accelerated implementation of the SMC method and apply it to investigate the nanostructure evolution in the $\rm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ and MoNbTaW HEAs, employing two simple machine learning energy models using the local short-range order parameters [48, 33, 49] as input features. We demonstrate the high efficiency of our method for large-scale MC simulation, which enables the simulation of atomistic systems exceeding one billion atoms using a single GPU (graphic processing unit). The large-scale simulation reveal a rich diversity of nanoscale phenomena, and we present a detailed analysis in the result section of this work.
@@ 2.1 Performance of the SMC algorithm
Our algorithm can be seen as a generalization of the checkerboard algorithm for the 2D Ising model, as explained in Supplementary 5.1. The key insight lies in rendering sequential Monte Carlo (MC) trials independent by partitioning the system into sufficiently large link-cells, thereby isolating the influence of individual MC trial moves. To simplify the explanation, a schematic of the link-cell for the 2D square lattice is shown in Fig. 1 (a). Each site is labeled by two indices $(i_{C},i_{A})$, where $i_{C}$ represent different link-cells and $i_{A}$ represents the different atoms within the link-cell. For the sake of discussion, let us assume that the atoms have only nearest-neighbor interactions. In such a case, the MC updates of the atoms with the same $i_{A}$ index but different $i_{C}$ index are independent of each other. This obviously presents a parallelism opportunity. For instance, consider atom $(1,13)$. During MC update, atom $(1,13)$ can swap with any of its nearest neighbors $(1,12)$, $(1,14)$, $(1,8)$ and $(1,18)$, which are colored as yellow in Fig. 1(a). Due to the short interaction range (nearest-neighbor), these moves would have no impact on calculating the energy changes for the swap moves of atom $(i_{C},13)$ (red sites in Fig. 1(a)), and the degree of parallelism is the number of cells $n_{C}$. We denote the yellow and green sites surrounding each red site in Fig. 1 as a local-interaction zone (LIZ), which is a name inspired by the locally self-consistent multiple scattering (LSMS) method [50, 51]. In other words, for any two sites, as long as the MC trial of one site does not affect the energies of the LIZ of the other site, then these two sites are independent. It is easy to see that the above discussion based on the 2D square lattice can be easily extended to the case of 3D crystal. Furthermore, a domain decomposition scheme is introduced in the SMC method to distribute the lattice across multiple GPUs, which further enhances the achievable system size, as illustrated in Fig. 1 (b). Generally, to determine the energy change resulting from the MC swap trial at site $i$, the local energies of each site within the local interaction zone (LIZ) of site $i$ must be evaluated. Additionally, chemical environment information extending beyond the LIZ is required for the energy change calculation, as illustrated by the LIZ+ region in Fig. 1 (c). Note that for a pair-interaction model, the above discussion can be simplified since the total energy changes can be calculated from the local energies of the two sites involved in the swap trial. As a result, the speed of the effective pair-interaction (EPI) model is faster than the generalized nonlinear model by approximately a factor of 50, as shown in Fig. 1 (d).
@@
Figure 1: Schematics of the scalable Monte Carlo (SMC) method. (a) Illustrate the SMC method with a 2D square lattice, in which each site is denoted by two indices $(i_{C},i_{A})$, where $i_{C}$ represents different link-cells, and $i_{A}$ represents the different atoms within the $n_{A}$ atom cell. (b) The whole chemical configuration is distributed on multiple GPUs, and the atoms near the boundary need to be communicated between GPUs $n_{A}$ times in every MC sweep. (c) A 2D view of the fcc or bcc lattice, site 1 (red) can swap with each of its nearest neighboring sits (yellow), and the green sites represents the ones that the local energies can be affected. The yellow and green sites, together with the centering atom form the LIZ. The chemical environment with LIZ+ are needed to calculate the energy change due to a swap trial. (d) A log-log plot of computation time vs the system size to illustrate the speed-up ratio of the SMC method as compared to DFT (MuST-KKR), linear-DFT (LSMS), and GNN (Allegro). The measured values for SMC are signified as stars, and ideal scaling assumed for all lines.
To illustrate the strength of our method, we compare the speed of the SMC method with other possible schemes to integrate MC with high-accuracy energy prediction methods, such as DFT, linear-DFT, and GNN (graph neural networks), as shown in Fig. 1, in which the lines are drawn by extrapolating from a data point using ideal scaling behavior. The actual computational speeds of the various methods are inherently influenced by numerous factors, such as the hardware, software implementations, material systems, and computational parameters. Therefore, the discussion presented here is intended solely to provide a theoretical estimation of their relative performance in terms of orders of magnitude. The stars in the figure represents the measured values of the SMC method, and the lines show the ideal scaling behaviors. For DFT and linear-DFT, we assume that an 100-atoms SCF calculation takes one hour, based on our experience with the MuST-KKR code [52]. The speed of the GNN method (Allegro) is estimated from the $\rm{Li_{3}PO_{4}}$ structure of 421,824 atoms in Ref. [24]. Note that a MC sweep is defined as making a MC trial move over each lattice sites. Combining it with the intrinsic $O(N^{3})$ scaling in the DFT method, the total computational cost of DFT then scales as $O(N^{4})$, as signified in Fig. 1 (d). Building on the principles of near-sightedness [53], the LSMS method employs an approximation that confines electron scattering to a local interaction zone. This approximation reduces the computational cost of energy evaluation to linear-scaling, therefor the total computational cost for a MC sweep scales approximately as $O(N^{2})$. For a one million-atom system, the decreased computational complexity can reduce the cost by a factor of $10^{8}$, as shown in the orange line in Fig. 1 (d). Nevertheles, linear-DFT still requires solving the Schrödinger equations explicitly. By replacing the computationally expensive DFT method with machine learning models such as the GNN model, the computationally cost can be further reduced by a factor of $10^{8}$, as shown in the green line in Fig. 1 (d). Moreover, our SMC parallelization strategy significantly reduces the computational cost by replacing the energy evaluation of the whole system with calculating the energy changes of the LIZ, which is independent of the system size. This strategy further decrease the computational complexity to $O(N)$, which amounts to a speedup of $10^{6}$ for a one million atom system. Finally, the SMC method can simultaneously harness the thousands or tens of thousands of cores in a modern high-performance GPU (e.g. 14,592 FP32 CUDA cores in an Nvidia H-800 GPU). For the EPI model, we estimate the GPU implementation contributes a speedup of $10^{3}$. Based on the preceding discussion, the total acceleration ratio of the SMC_GPU method relative to a CPU based DFT method, for a system of $N$ atoms, can be decomposed into contributions from nearsightedness $S_{NS}$, ML acceleration $S_{ML}$, SMC method $S_{SMC}$, and GPU acceleration $S_{GPU}$, and written as:
$S=S_{NS}\times S_{ML}\times S_{SMC}\times S_{GPU}\approx(N/100)^{2}\times 10^{8}\times N\times 10^{3}=10^{7}\times N^{3},$ (1)
@@
as illustrated in Fig. 1 (d). Therefore, for a one million atom system, we have $S=10^{25}$, as highlighted in Fig. 1 (d). We can also see that for a nonlinear model, such as the Local_SRO model introduced in the following section, the speedup will be reduced by a factor of 50, due to the aforementioned necessity of evaluating the local energies of each sites within the LIZ. If we increase the system size to one billion atoms, then the speedup of SMC with respect to DFT will be further increased by a factor of $10^{9}$ to reach $10^{34}$, as highlighted in Fig. 1 (d). This analysis clearly demonstrates the exceptional performance of the SMC method for large-scale MC simulations.
@@
In addition to theoretical estimations, we compare our work with selected prior studies on the atomistic simulation of high-entropy alloy (HEA) thermodynamics, as summarized in Tab. 1. To contextualize the simulation scale of our work, we extend our search beyond HEAs to identify the largest atomistic simulation system reported in the literature. A recent study achieved this with a system of 29 billion atoms, which employs 35 million CPU cores on one of the world’s largest supercomputers (comprising 90,000 computing nodes, approximately 84% of the entire system). In contrast, constrained by computational resources, our work utilized only two Nvidia H-800 GPUs, yet achieved a simulation scale of one billion atoms. Note that this already surpasses the 100-million-atom system size of the 2020 Gordon-Bell Prize winner [22]. For HEAs, our work represents the largest system size achieved in atomistic simulations at ab initio accuracy. Furthermore, we demonstrate that, through careful design of the Monte Carlo (MC) algorithm to fully exploit its inherent parallelization potential, MC simulations can be scaled to billions of atoms, rivaling the scalability of molecular dynamics (MD) methods, which are traditionally viewed as more amenable to parallelization and large-scale simulations. Second, we highlight that, due to limitations in accessible system sizes, no prior work has addressed the challenge of modeling nanostructures in HEAs, particularly when the feature size of these nanostructures requires millions of atoms for accurate description. Our capability not only underscores the strength of our method but also addresses a critical gap in the field, as nanoparticles[40, 47, 54, 39] and nonophases [55] play a pivotal role in the design of HEAs with superb mechanical properties, such as overcoming the trade-off in strength and ductility [55, 39].
@@
Table 1: Comparison of our work with selected prior studies employing atomistic simulations in high-entropy alloys. For reference, we also include the largest atomistic system (29-B) simulated using machine learning (ML) models, as identified in the literature. Notably, Ref. [59] utilized 35 million CPU cores in one of the world’s fastest supercomputers, whereas this work achieved billion atoms using only two Nvidia H-800 GPUs.

### v2

@@ Abstract
The computational cost of traditional first-principles method quickly becomes prohibitively expensive as the number of atoms increases. This challenge is further amplified by the need to evaluate finite-temperature properties with Monte Carlo (MC) simulations, which is inherently challenging to parallelize due to sequential Markov chain updates. Here, we introduce Scalable Monte Carlo (SMC), an efficient MC simulation method that overcomes the parallelization bottlenecks in conventional MC simulation, reducing the computational complexity of a MC sweep from quadratic to linear. We present a GPU implementation of the SMC method, SMC-GPU, which simultaneously harnesses the thousands of processing cores on a GPU to accelerate the computation. By adopting a data-driven workflow that surrogates the computationally expensive density functional theory (DFT) with ML models, we demonstrate that SMC-GPU is capable of simulating systems of more than one-billion atoms, while maintaining the accuracy of first-principles methods. Using this unprecedented capability, we performed billion-atom MC simulations to investigate the nanostructure evolution of two important high-entropy alloys (HEAs), FeCoNiAlTi and MoNbTaW, in which the NPs are believed to be responsible for their superb mechanical properties. Our results reveal a rich diversity of nanostructures, including nanoparticles (NP), 3D-connected NP, and disorder protected nanophases. We quantitatively analyze the size, composition, and morphology of the nanostructures, as well as directly simulate the atom-probe-tomography (APT) needle. The results align well with available experimental observations. This work underscores the promising potential of leveraging large-scale MC simulation to explore the largely uncharted territory of nanostructure evolution in HEAs.
@@ 1 Introduction
A central goal of computational materials science is to predict the physical properties of materials using only fundamental inputs such as atomic species and physical constants. A widely adopted strategy towards this goal is the direct solution of the quantum equations governing electron dynamics: the Schrödinger equation for non-relativistic cases and the Dirac equation for relativistic cases. However, the complexity of quantum many-body interactions necessitates approximations. For such a purpose, density functional theory (DFT) [1] has emerged as a particularly successful approach, which approximates intractable many-body effects through an exchange-correlation function within a single-body framework. This approximation significantly enhances computational efficiency, enabling the routine prediction of ground-state physical properties for systems with several hundred atoms. However, for systems requiring more than thousands of atoms to simulate, such as nanodefects, non-stoichiometric compounds, and complex solid-solution alloys, conventional DFT methods become prohibitively expensive due to their intrinsic $O(N^{3})$ scaling behavior with system size. A further complication arises when considering finite-temperature effects [2]. Even for relatively small systems, rigorously accounting for temperature via ensemble averaging in statistical mechanics necessitates the evaluation of an enormous number of configurations, making DFT-based finite-temperature simulations prohibitively expensive. For example, a direct thermodynamic simulation of the CuZn alloy using a 250-atom supercell requires the calculation of 600,000 DFT energies [3], a task demanding some of the world’s largest supercomputers.
@@
One common strategy to accelerate atomistic simulations is to replace computationally expensive electronic structure calculations with efficient atomistic models. From the perspective of renormalization group theory [4, 5], the atomistic models can be interpreted as the effective model obtained by integrating out the electronic degrees of freedom, which substantially reduces the computational cost. Indeed, traditional empirical models, such as the embedded atom model (EAM) and cluster expansion methods, have been widely employed for large-scale thermodynamic simulations. However, traditional empirical models are limited in terms of functional form, which restricts their generalization capability and predictive accuracy. On the other hand, the advent of machine learning (ML) for atomistic systems [6, 7, 8, 9] has introduced a new paradigm to the construction of effective atomistic models [10, 11, 12, 13, 14]. Instead of being predefined, the parameters in ML atomistic models are automatically determined by training from high-quality first-principles datasets, therefore rendering it possible for the model to automatically capture complex interatomic interactions [15, 16, 17, 18]. Well-trained ML models are typically orders of magnitude faster than DFT methods, while can still retain their high accuracy [19, 20, 21]. This renders it possible to carry out high-accuracy atomistic simulations using supercells containing millions of atoms [22, 23, 24, 25], far exceeding the thousands-of-atoms limits of conventional DFT approaches.
@@
However, most machine learning potentials have been developed primarily to accelerate molecular dynamics (MD) simulations [22, 23, 7, 26, 27], with relatively limited applications in Monte Carlo (MC) simulations [28, 29, 16, 30, 29]. This is notable given that MC represents one of the two cornerstone methods in atomistic simulations, alongside MD [31]. A key challenge lies in the intrinsic sequential updating nature of widely used MC algorithms, such as the Metropolis algorithm. The MC trials are generally attempted site by site, which hinders large-scale parallelization. This is in stark contrast to MD, where all the atoms are updated simultaneously in a single step. The sequential updating nature of MC quickly becomes the bottleneck as the system size grows, which undermines the efficiency advantage of ML atomistic models when integrating the two. On the other hand, for many interesting phenomena [32, 33, 34], such as order-disorder transitions in chemically complex materials, MC simulations remain the only viable approach due to their sampling efficiency. Consequently, developing highly scalable MC algorithms that overcome the limitations of sequential updating is of paramount importance for realizing the potential of ML atomistic models.
@@
To overcome the parallelization bottleneck inherent in conventional MC simulations, we introduce the Scalable Monte Carlo (SMC) method, which is a generalized checkerboard algorithm [35] that reduces the computational complexity of an MC sweep from $O(N^{2})$ to $O(N)$. Compared to the checkerboard algorithm variant [35] that only applies to Ising interactions, we demonstrate that the SMC method can be generalized to arbitrary local interaction models by introducing a local interaction zone (LIZ), which makes it suitable for integration with machine learning atomistic models. We employ our method to study high-entropy alloys (HEAs) [36, 37], a class of chemically complex materials that have received significant attention due to their exceptional mechanical properties. These properties include overcoming the traditional strength-ductility trade-off [38, 39, 40], attributed to phenomena such as chemical short-range order [41, 42, 43], nanoprecipitates [39], and nanophases [44]. Understanding these features necessitates large-scale MC simulations. Beyond mechanical properties, the size and morphology of nanostructures in HEAs also offer promising opportunities for catalysis, sparking widespread interest in recent years [45, 46, 47]. We present a GPU-accelerated implementation of the SMC method, SMC-GPU, and apply it to investigate the nanostructure evolution in the $\rm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ and MoNbTaW HEAs, employing two simple machine learning energy models using the local short-range order parameters [48, 33, 49] as input features. We demonstrate the excellent efficiency and scalability of the SMC method, which enables the simulation of atomistic systems exceeding one billion atoms using a single GPU (graphic processing unit). Our results reveal a rich diversity of interesting nanoscale phenomena, including nanoparticles (NP), 3D-connected NP network, and disorder protected nanophases. We quantitatively analyze the size, composition, and morphology of the nanostructures, which align well with available experimental results obtained with atom-probe-tomography (APT) and electron microscopy. Finally, our results reveal that the intricate nanoscale interplay of order and disorder in high-entropy alloys (HEAs) stems from the combined effects of chemical complexity and temperature, offering valuable guidance for alloy design.
@@ 2.1 Performance of the SMC algorithm
The SMC algorithm can be seen as a generalization of the checkerboard algorithm for the 2D Ising model, as explained in Method 3.2 and Supplementary 5.1. The key insight lies in rendering sequential Monte Carlo (MC) trials independent by partitioning the system into sufficiently large link-cells, thereby isolating the influence of individual MC trial moves. To simplify the explanation, a schematic of the link-cell for the 2D square lattice is shown in Fig. 1 (a). Each site is labeled by two indices $(i_{C},i_{A})$, where $i_{C}$ represents different link-cells and $i_{A}$ represents the different atoms within the link-cell. For the sake of discussion, let us assume that the atoms have only nearest-neighbor interactions. In such a case, the MC updates of the atoms with the same $i_{A}$ index but different $i_{C}$ indexes are independent of each other. This obviously presents a parallelism opportunity. For instance, consider atom $(1,13)$. During MC update, atom $(1,13)$ can swap with any of its nearest neighbors $(1,12)$, $(1,14)$, $(1,8)$ and $(1,18)$, which are colored as yellow in Fig. 1(a). Due to the short interaction range (nearest-neighbor), these moves would have no impact on calculating the energy changes for the swap moves of atom $(i_{C},13)$ (red sites in Fig. 1(a)), and the degree of parallelism is the number of cells $n_{C}$. We denote the yellow and green sites surrounding each red site in Fig. 1 as a local-interaction zone (LIZ), which is a name inspired by the locally self-consistent multiple scattering (LSMS) method [50, 51]. In other words, for any two sites, as long as the MC trial of one site does not affect the energies of the LIZ of the other site, then these two sites are independent. It is easy to see that the above discussion based on the 2D square lattice can be extended to the case of 3D crystal. Furthermore, in the SMC method, a domain decomposition scheme is introduced to distribute the lattice among multiple GPUs, which further enhances the achievable system size, as illustrated in Fig. 1 (b). Generally, to determine the energy change resulting from the MC swap trial at site $i$, the local energies of each site within the local interaction zone of site $i$ must be evaluated. Additionally, chemical environment information extending beyond the LIZ is required for the energy change calculation, as illustrated by the LIZ+ region in Fig. 1 (c). Note that for a pair-interaction model, the above discussion can be simplified since the total energy changes can be calculated from the local energies of the two sites involved in the swap trial. As a result, the speed of the effective pair-interaction (EPI) model is faster than the generalized nonlinear model by approximately a factor of 50, as shown in Fig. 1 (d).
@@
Figure 1: Schematics of the scalable Monte Carlo (SMC) method. (a) Illustrate the SMC method with a 2D square lattice, in which each site is denoted by two indices $(i_{C},i_{A})$, where $i_{C}$ represents different link-cells, and $i_{A}$ represents the different atoms within the $n_{A}$ atom cell. (b) The whole chemical configuration is distributed on multiple GPUs, and the atoms near the boundary need to be communicated between GPUs $n_{A}$ times in every MC sweep. (c) A 2D view of the fcc or bcc lattice, site 1 (red) can swap with each of its nearest neighboring sites (yellow), and the green sites represent the ones that the local energies can be affected. The yellow and green sites, together with the centering atom form the LIZ. The chemical environment with LIZ+ are needed to calculate the energy change due to a swap trial. (d) A log-log plot of computation time vs the system size to illustrate the speed-up ratio of the SMC method as compared to DFT (MuST-KKR), linear-DFT (LSMS), and GNN (Allegro). The measured values for SMC are signified as stars and ideal scaling is assumed for all lines.
@@
To illustrate the strength of our method, we compare the theoretical speed of the SMC method with other schemes to integrate MC with high-accuracy energy prediction methods, such as DFT, linear-DFT, and GNN (graph neural networks), as shown in Fig. 1 (d). Note that the lines are drawn by extrapolating from a data point using ideal scaling behavior. The actual computational speeds of the various methods are inherently influenced by many factors, such as hardware, software implementations, material systems, and computational parameters. Therefore, the discussion presented here is intended solely to provide a theoretical estimation of their relative performance in terms of orders of magnitude. The stars in the figure represent the measured values of the SMC method, and the lines show the ideal scaling behaviors, which are explained in Method 3.1. From Fig. 1 (d), we see that for a one million atom system, we have the speedup $S=10^{25}$, as highlighted in Fig. 1 (d). We can also see that for a nonlinear model, such as the Local_SRO model introduced in the following section, the speedup will be reduced by a factor of 50, due to the aforementioned necessity of evaluating the local energies of each site within the LIZ. If we increase the system size to one billion atoms, then the speedup of SMC with respect to DFT will be further increased by a factor of $10^{9}$ to reach $10^{34}$, as highlighted in Fig. 1 (d). This analysis clearly demonstrates the exceptional performance of the SMC method for large-scale MC simulations.
@@
In addition to theoretical estimations, we compare our work with selected previous studies, as summarized in Tab. 1. It can be seen that our work represents the largest system size achieved in atomistic simulations at ab initio accuracy. Such capability enables us to directly study large nanostructures comprising of millions of atoms, in contrast to previous works that are limited to short-range order (SRO), long-range order (LRO), and small nanoprecipitates. To contextualize the simulation scale of our work, we extend our search beyond HEAs to identify the largest atomistic simulation system reported in the literature. A recent study achieved this with a system of 29 billion atoms, which employs 35 million CPU cores on one of the world’s largest supercomputers (comprising 90,000 computing nodes, approximately 84% of the entire system). In contrast, constrained by computational resources, our work utilized only two NVIDIA H800 GPUs, yet already achieved a simulation scale of one billion atoms. Note that this already surpasses the 100-million-atom system size of the 2020 Gordon-Bell Prize winner [22]. Furthermore, the comparison presented above demonstrates that, through careful design of the Monte Carlo algorithm to fully exploit its inherent parallelization potential, MC simulations can rival molecular dynamics (MD) methods in terms of scalability, which are traditionally viewed as advantageous over MC methods in simulation scale. Finally, we highlight that the excellent computational efficiency of the SMC method not only presents a success in pushing the simulation length scale forward, but also provides a particularly important tool for understanding the vital role of nanostructures [52, 40] in the exceptional mechanical properties in HEAs [47, 53, 39], such as overcoming the trade-off in strength and ductility [52, 39].
@@
MoNbTaW Hybrid MC/MD SNAP 580,000 Dislocation motion under SRO and nanoscale B2 precipitates [57]
@@
Table 1: Comparison of our work with previous studies employing atomistic simulations for high entropy alloys. For reference, we also include the largest atomistic system (29-B) simulated using machine learning (ML) models, as identified in the literature. Notably, Ref. [58] utilized 35 million CPU cores in one of the world’s fastest supercomputers, whereas this work achieved a billion atoms using only two NVIDIA H800 GPUs.